# Score Operator Newton Transport

### 08 Dec 2023 v2

@@ 6 Discussion
Learning elliptic PDEs: Many structure-exploiting, fast, and sample-efficient methods are emerging for learning the solution operators of linear elliptic PDEs; see, e.g., Lu et al. [2021a], Boullé and Townsend [2023], Schäfer and Owhadi [2021]. These methods use randomized numerical linear algebra [Boullé and Townsend, 2023], CNN-based encoder-decoder networks [Zhang and Garikipati, 2023], interpolation between deep neural network and Monte Carlo approximations [Nüsken and Richter, 2023], etc. These results suggest that it is possible to develop optimal methods, in terms of computational and sample complexity, for learning our SCONE update operator $\mathcal{L}(q)^{-1}$ by exploiting low-rank structure in its solutions. With the same goal, we will also explore particle methods (e.g., from fluid dynamics [Monaghan, 2012, Cottet et al., 2000]), which can also use fast numerical linear algebra and have theoretical guarantees.
@@ Appendix C Additional experiments
fLS4ZhTiRJgqIoVBRFIoric47jPnmeB1mW/9rr9ZpSSn3Lsmir1fJZlqWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 09 Mar 2024 v3

@@ 6 Discussion
Learning elliptic PDEs: Many structure-exploiting, fast, and sample-efficient methods are emerging for learning the solution operators of linear elliptic PDEs; see, e.g., Lu et al. [2021a], Boullé and Townsend [2023], Schäfer and Owhadi [2021]. These methods use randomized numerical linear algebra [Boullé and Townsend, 2023], CNN-based encoder-decoder networks [Zhang and Garikipati, 2023], interpolation between deep neural network and Monte Carlo approximations [Nüsken and Richter, 2023], etc. These results suggest that it is possible to develop optimal methods, in terms of computational and sample complexity, for learning our SCONE update operator by exploiting low-rank structure in its solutions. With the same goal, we will also explore particle methods (e.g., from fluid dynamics [Monaghan, 2012, Cottet et al., 2000]), which can also use fast numerical linear algebra and have theoretical guarantees.
@@
Funding NC and YMM acknowledge support from NSF grant PHY-2028125. YMM acknowledges support from DOE ASCR award DE-SC0023187, AFOSR award FA9550-20-1-0397, and ONR award N00014-20-1-2595. FS acknowledges support from AFOSR award FA9550-23-1-0668 and ONR award N00014-23-1-2545.
Acknowledgments We are greatly indebted to Guillaume Bal for kindly pointing out that the operator $\mathcal{L}$ is not elliptic, as we had assumed in our earlier draft, and generously helping us correct the error. We would also like to thank Daniel Sharp for providing us with an mParT implementation, which we could readily use to generate Figure 4.